hello.sir..plzz reply as soon as possible..i have an test on sunday 8 aug...


..there  is an arc OAB  of radius 100m.with centre O.. which substend angle of  90 degree at the centre .a car of wieght 100 kg moves on arc.


what will be  the friction force at point A and mid piont of arc.??


speed of vehicle is 5 m per s.


 


thankxxx

Grade:12

1 Answers

Gokul Joshi AskiitiansExpert-IITK
13 years ago

Dear Divyavikram,

In any curved segment the acceleration is v2/r (no tangential acceleration because going with constant speed) so the force acting must be mv2/r = 100*52/100 = 25 N. In any linear segment acceleration is zero so force needed is zero.

Please feel free to post as many doubts on our discussion forum as you can.we will get you the answer and detailed  solution very  quickly. All the best.

Regards,
Askiitians Experts,
Gokul  Joshi